OpenStudy (anonymous):

Which of the following shows the right side of the equation after applying a reasonable first step to solve negative h divided by two plus negative one-fourth equals four-ninths? four-ninths minus one-fourth four-ninths plus four-ninths four-ninths plus one-fourth four-ninths minus four-ninths

OpenStudy (anonymous):

@johnweldon1993 can you help me please

OpenStudy (johnweldon1993):

\[-\frac{h}{2} - \frac{1}{4} = \frac{4}{9}\] What you really want to do is isolate that -h/2 part first...so add 1/4 to both sides of the equation... so the right side would look like... \[\large \frac{4}{9} + \frac{1}{4}\] Which is answer choice...?